Circular No. 8875

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007gw–2007kd Further to IAUC 8874, published data for objects not yet spectroscop-

ically confirmed appear on the specified CBETs: 2007gw in NGC 4161, CBET 1042 (Corrigendum to CBET 1042: 2006dk was a type-II supernova); 2007ig in MCG –01-5-2, CBET 1060; 2007iq in UGC 3416, CBET 1064; 2007iw in ESO 325-43, CBET 1071; 2007kc in UGC 8400, CBET 1082; 2007kd in MCG +6-21-36, CBET 1083. The spectral type of 2007ir in UGC 2033 is unknown (cf. CBETs 1067, 1077). Data for type-Ia supernovae: 2007hj in NGC 7461, CBET 1048; 2007hu in NGC 6261, CBETs 1056, 1058; 2007if, CBET 1059; 2007is in UGC 10553, CBETs 1067, 1075. SN 2007hb in NGC 819 is a type-Ib/c supernova (cf. CBETs 1043, 1066, 1067). Data for type-II supernovae: 2007hv in UGC 2858 (CBETs 1056 and 1058); 2007il in IC 1704, CBET 1062; 2007it in NGC 5530, CBETs 1065, 1068; 2007iv in UGC 12917, CBETs 1070, 1072. Prediscovery V magnitudes of 2007it from G. Pojmanski (Warsaw University Astronomical Observatory) from All Sky Automated Survey images, somewhat affected by light from the galactic core: Sept. 4.008 UT, [14; 5.994, 12.4; 6.014, 12.0; 8.036, 12.4; 11.027, 12.1; 12.983, 12.3. SNe 2007gx, 2007hl, and 2007hn are also of type Ib or Ic; 2007gy, 2007gz, 2007hc, 2007hi, 2007hm, 2007hr, 2007hs, 2007hw, 2007ib, 2007iu, 2007iz, 2007ja, 2007jf, 2007jm, and 2007jn are of also of type II; the remainder are type-Ia supernovae. SN 2007iu is 10′′.3 west and 1′′.3 north of the center of the apparent host galaxy (which is not a star, as stated on CBET 1069). Objects questioned as possible type-Ia supernovae include 2007ie, 2007ii, 2007jc, and 2007jv. Objects questioned as possible type-Ib or -Ic supernovae include 2007jc, 2007ju, and 2007jy.

SUPERNOVAE 2007ke, 2007kf, 2007kg, 2007kh Four additional apparent supernovae have been found on unfiltered

CCD images: 2007ke in NGC 1129 by J. Chu and W. Li (LOSS/KAIT; details on CBET 1084); 2007kf by W. Kloehr (cf. IAUC 8851; details on CBET 1085); 2007kg by H. Mikuž (cf. IAUC 8834; details on CBET 1086); and type-Ia SN 2007kh by F. Yuan et al. (details on CBET 1089). Discovery observations tabulated for those objects reported to be around mag 17:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007kf Sept. 22.93 17^{h}31^{m}31.^{s}76 +69^{o}18′40′′.1 17.0 2′′.2 E, 4′′.9 S
2007kg Sept. 22.90 23 58 37.43 +60 59 06.7 17.1 7′′ E, 6′′ N
2007 September 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green